OpenStudy (anonymous):

help asap !! ill give medals The table below represents a linear function f(x) and the equation represents a function g(x). x f(x) 0 30 1 32 2 34 g(x) g(x) = 8x - 25 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (anonymous):

equation for the table is y = 2x + 30 slope (32 -30)/(1-0) = 2 when x = 0, y = 30 linear equation the slope intercept form is y = mx + b m is the slope of the line b is the y-intercept, or where the line cuts the y axis so f(x) has a slope of 2 g(x) has a slope of 8 the larger the number for the slope, the steeper the line.

OpenStudy (anonymous):

question 1. you can say, the point slope form of a line is y = mx + b. f(x) has a slope of 2 and g(x) has a slope of 8, g(x) is the line with the steepest slope. Question 2 f(x) has a y intercept of 30 and g(x) has a y intercept of -25... so f(x) has the greater intercept. hope this helps